Circular No. 9286

Central Bureau for Astronomical Telegrams

Postal Address: Hoffman Lab 209; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbatiau@eps.harvard.edu  ISSN 0081-0304  cbat@iau.org

               URL http://www.cbat.eps.harvard.edu/index.html

Prepared using the Tamkin Foundation Computer Network

V5856 SAGITTARII = PNV J18205200–2822100 As first announced on CBET 4334, S. Nakano reported the discovery of

an apparent nova (mag 10.4) by Yukio Sakurai (Mito, Ibaraki-ken, Japan) on two 10-s digital exposures taken on 2016 Oct. 26.38 UT with a 180-mm-focal-length f/2.8 camera lens (+ Nikon D7100 camera); Sakurai measured the following position for the variable: α = 18^{h}20^{m}52^{s}.0, δ = −28^{o}22′10′′ (equinox 2000.0). He adds that nothing is visible at this position on two frames taken on Oct. 23.398 using the same instrumentation. The variable was designated PNV J18205200–2822100 when it was posted at the Central Bureau’s TOCP webpage. Nakano measured Sakurai’s jpeg image and found mag 10.7 and position end figures 52^{s}.42, 11′′.2 (± 2′′), adding that the limiting magnitude seems to be 12.5. P. Schmeer, Bischmisheim, Germany, noted that an independent discovery of PNV J18205200–2822100 was made by K. Z. Stanek et al. in the course of the All-Sky Automated Survey for Supernova (ASAS-SN) using the robotic 14-cm “Cassius” telescope at Cerro Tololo, in which the variable appeared at magnitude V ∼ 13.7 on Oct. 25.02 (with position end figures 52^{s}.12, 13′′.5); additional ASAS-SN magnitudes for PNV J18205200–2822100: Oct. 20.04, [17.3; 26.02, 11.6.

P. Luckas, University of Western Australia, obtained low-resolution CCD spectroscopy (range 380–720 nm) of PNV J18205200–2822100 on 2016 Oct. 27.5 UT with a 35-cm reflector at Shenton Park Observatory, Perth. The spectrum suggests an optically thick stage of nova expansion; the Balmer series were seen in emission through Hδ with little evidence for any P-Cyg absorption, with a low maximum radial velocity (HWZI) of −750 ± 50 km/s at Hα and Hβ. There is, however, a broader weak extended wing at Hα (at ∼ 5 percent of the continuum and only a few percent of the peak intensity) that extends to nearly 1050 km/s.

Further details are given on CBET 4334. N. Samus and E. Kazarovets note that this variable has been given the permanent GCVS designation V5856 Sgr. This is presumably also Nova Sgr 2016 No. 4 (cf. IAUC 9284).

COMET C/2017 E2 (TSUCHINSHAN) The IAU Working Group on Small Body Nomenclature has voted to

change the name of this comet (cf. CBET 4370) from “XuYi” to “Tsuchinshan”, following the pattern established with comets 62P/1965 A1 and 60P/1965 A2 (cf. IAUC 1892).

COMET C/2017 E4 (LOVEJOY) Visual total-magnitude estimates by J. J. Gonzalez, Leon, Spain, with

10×50 binoculars: Mar. 28.18 UT, 7.9; Apr. 1.17, 7.4; 7.17, 7.1.

2017 April 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2017 CBAT  Daniel W. E. Green